OpenStudy (anonymous):

Express a given rational number as a quotient of two integers: 0.1111111....

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@JeanneChristine Are you still there?

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (anonymous):

let a = 0.11111.... 10a = 1.111111...... 10a-a = 1 9a = 1 a=1/9

OpenStudy (anonymous):

@rizwan_uet an elegant answer, but you should guide the asker to the answer instead of giving it to them.

OpenStudy (anonymous):

I think technically speaking that if you have 0.1111 . . . (on and on) it's really 0.2

OpenStudy (anonymous):

I think for the very reason that rizwan_uet solved the problem.

OpenStudy (anonymous):

skullpatrol, yeah. When I was on here yesterday some kid was asking for the answer to a test question, while he was taking it. o.O

OpenStudy (anonymous):

Peter14, I dunno about this particular question. I wouldn't have been able to be guided to the answer without seeing the answer that rizwan_uet gave.

OpenStudy (anonymous):

suppose 2.02123123123................ we see that number 123 is repeated again and again. Multiply R.H.S again and again by 10 till we get again123 .Here i do Let x=2.02123123123....... ....(1) 10x=20.2123123123.......... .....(2) 100x=202.123123123....... .....(3) 1000x=2021.23123123...... .....(4) 10000x=20212.3123123...... .....(5) 100000x=202123.123123...... ....(6) we see in (3) and (6) part after the decimal is same. So subtract (3) from (6) (100000-100)x=202123.123123......-202.123123... 99900x=201921 \[x=\frac{ 201921 }{99900 }\]